7. A 24.6 ml solution of unknown acid contains 0.0034 mols of acid (5.03 g/L) in solution.
Calculate the molar mass of the solution, and use that to calculate the concentration and
pH of the solution.
8.) Which of the 4 acids was the acid from problem 7?
a) H2SO4
b) HCl
c) HNO3
d) HI
1
Expert's answer
2020-05-10T14:57:46-0400

Solution:

The molar mass of unknown acid = (Density of solution × Volume of solution) / (Moles of acid).

Units: [g/mol] = (g/L × L) / (mol) = [g/mol] (Correct).

Then,

The molar mass of unknown acid = (5.03 g/L × 0.0246 L) / (0.0034 mol) = 36.39 g/mol


The molar mass of H2SO4 is 98.079 g/mol.

The molar mass of HCl is 36.46 g/mol.

The molar mass of HNO3 is 63.01 g/mol.

The molar mass of HI is 127.911 g/mol.

The molar mass of unknown acid = 36.39 g/mol.

Therefore, the unknown acid is HCl.


The mass of solution = Density of solution × Volume of solution

The mass of solution = (5.03 g/L) × (0.0246 L) = 0.1237 g


Molarity of HCl = Moles of HCl / Volume of solution

Molarity of HCl = (0.0034 mol) / (0.0246 L) = 0.1382 mol/L

Molarity of HCl = 0.1382 mol/L


Hydrochloric acid, HCl, is a strong acid that dissociates completely in water to form H3O+ and Cl.

HCl(aq) + H2O(l) ⇄ H3O+(aq) + Cl(aq).

Molarity of HCl = CM(HCl) = [H3O+] = 0.1382 mol/L.

[H3O+] = 0.1382 M


The pH is calculated using the expression:

pH = - log[H3O+]

pH = - log(0.1382) = 0.86

pH = 0.86


Answer (7):

The molar mass of unknown acid = 36.39 g/mol

Molarity of HCl = 0.1382 M

pH = 0.86


Answer (8):

The unknown acid is HCl.